Let u=8i - 8j, and w=-i-2j. Find ||w-ul. w-ul=1 (Type an exact answer, using radicals as needed.)

37


Let u=8i - 8j, and w=-i-2j. Find ||w-ul. ||w-ul=1 (Type an exact answer, using radicals as needed.)
Answer #1

U = 01 - 89 W = -1 --1-2 I woull 11 W - ull Jawoue w-u> 7-99 +6 + 6j1 won = (-î -aſ) -181-89) (-gî +69) (wou, wous -qî x 6 7